I have alot of questions similar to this, and I would like an example done for me. I don't understand the Law of Cosines very well... heres one question: If a=4, b=5, and c=7, find angle A. thanks!
Answer by stanbon(48564)

Law of Cosines: a^2=b^2+c^2-2bc(Cos A)
Solving for Cos A you get :
Cos A = [b^2+c^2-a^2]/2bc
In your problem you get
Cos A = [5^2+7^2-4^2]/2(5)(7)
Cos A = 58/70
Cos A = 0.8286
Find the inverse Cosine of 0.8286
A=34.05 degrees
Cheers,
stan H.